slove the following system of equations
x+5y=8
x+y=4

Found 3 solutions by richwmiller, stanbon, blwinbbbles:
Answer by richwmiller(17219)

x+5y=8
x+y=4
subtract the second from the first
4y=4
y=1
x=3

Answer by stanbon(75887)

solve the following system of equations
x+5y=8
x+y=4
-------
Subtract the 2nd equation from the 1st; then solve for "y":
4y = 4
y = 1
------------
Substitute y=1 into x+y = 4 to solve for "x":
x + 1 = 4
x = 3
------------------
Cheers,
Stan H.

Answer by blwinbbbles(106)

you have
(1)x+5y=8 and
(2)x+y=4
you can use elimination:
Multiple equation (2) by -1
so now you have:
(1) x+5y=8
(2) -x-y=-4
add the equations and now you have
4y=4 divide both sides by 4
y=1
You now know y so sub into original equation (1)
x + 5(1) = 8
x+5=8 subtract 5 from both sides
x=3
your answer is (3,1)